OpenStudy (anonymous):

A) If \[x_1, ..., x_n\]are distinct numbers, find a polynomial function \[f_i\]of degree \[n-1\] which is 1 at \[x_i\] and 0 at \[x_j\] for \[j \neq i\] B) Now find a polynomial function \[f\] of degree \[n-1\] such that \[f(x_i) = a_i\]where \[a_1, ..., a_n\]are given numbers. You should use the functions \[f_i\] from part a

OpenStudy (asnaseer):

I think solution to A is:\[f_i=(x-x_1)(x-x_2)...(x-x_i+1)(x-x_{i+1})...(x-x_n)\]

OpenStudy (asnaseer):

sorry:\[f_i=\frac{(x-x_1)(x-x_2)...(x-x_i+1)(x-x_{i+1})...(x-x_n)}{(x_i-x_1)(x_i-x_2)...(x_i-x_j)...(x_i-x_n)}\]where no x_j term in the denominator is equal to x_i

OpenStudy (asnaseer):

then B would just be:\[f(x_i)=\frac{(x-a_1)(x-a_2)...(x)(x-a_{i+1})...(x-a_n)}{(x_i-a_1)(x_i-a_2)...(a_i-a_j)...(a_i-x_n)}\]

OpenStudy (zarkon):

wouldn't (B) be \[f(x)=\sum_{i=1}^{n}a_{i}f_{i}(x)\]

OpenStudy (asnaseer):

of course - now why didn't I spot that |-)